# Learning to Optimize for Mixed-Integer Non-linear Programming

### v2

@@ 5 Learning to Optimize MINLPs with Correction Layers
The two baselines involve only some parts of our full pipeline and thus serve as ablations. Rounding after Learning (RL) is used as a baseline, applying rounding to the relaxed solution generated by learning-based methods rather than learning-to-round as our correction layers do. In Rounding with STE (RS), the continuous values predicted by the first neural network $\pi_{\Theta_{1}}$ are directly rounded without any learned correction mechanism, i.e., without using the correction network $\varphi_{\Theta_{2}}$; STE is used to backpropagate through the rounding operator. For further details, refer to Appendix C.
@@ 6.3 Multi-Dimensional Mixed-Integer Rosenbrock Problem
where $\bm{x}\in\mathbb{R}^{n}$ are continuous decision variables and $\bm{y}\in\mathbb{Z}^{n}$ are integer decision variables. The vectors $\bm{p}\in\mathbb{R}^{n}$ and $\bm{q}\in\mathbb{R}^{n}$ are fixed for each instance, while the parameters $b$ and $\bm{a}$ vary. Details on how the parameters $\bm{p},\bm{q},\bm{a},$ and $b$ were generated can be found in Section D.2.
@@
RC Obj Mean $23.30$ $56.99$ $505.47$ $6245.15$ $70098.23$
Obj Median $21.22$ $50.96$ $439.79$ $6500.15$ $68253.48$
% Infeasible $2\%$ $1\%$ $1\%$ $2\%$ $32\%$
@@
6.4 Effect of Increasing Penalty Parameters
In this section, we evaluate the effect of progressively increasing penalty parameters on the above benchmark problems. This experiment is inspired by penalty methods commonly used in optimization, such as the Augmented Lagrangian Method (Hestenes, 1969). The weight of the penalty term is typically increased throughout the iterations. This gradual increase in penalty helps avoid numerical difficulties caused by overly strict constraints at the early stages while enhancing constraint satisfaction as the solution converges.
Table 5: Results for Increasing Penalty Parameters on Benchmark Problems. See the caption of Table 2 for details.
Problem Size Method Obj Mean Obj Median % Infeasible Time (Sec)
Convex Quadratic 100×100 RC $-10.660$ $-10.697$ $1\%$ $0.0036$
LT $-11.548$ $-11.617$ $0\%$ $0.0036$
200×200 RC $-25.350$ $-25.412$ $0\%$ $0.0033$
LT $-25.110$ $-25.297$ $0\%$ $0.0033$
500×500 RC $-60.040$ $-60.405$ $4\%$ $0.0062$
LT $-57.507$ $-57.590$ $1\%$ $0.0063$
Simple Non-Convex 100×100 RC $1.773$ $1.718$ $0\%$ $0.0038$
LT $1.932$ $1.917$ $0\%$ $0.0035$
200×200 RC $1.662$ $1.584$ $0\%$ $0.0037$
LT $2.163$ $2.060$ $0\%$ $0.0034$
@@ 6.4 Effect of Increasing Penalty Parameters
Rosenbrock 200×5 RC $519.85$ $467.59$ $0\%$ $0.0034$
LT $618.51$ $624.10$ $0\%$ $0.0031$
2000×5 RC $7276.96$ $7460.26$ $0\%$ $0.0042$
LT $5780.85$ $5590.29$ $0\%$ $0.0044$
20000×5 RC $63626.35$ $58977.61$ $29\%$ $0.0044$
LT $63734.97$ $64431.54$ $39\%$ $0.0044$
In this experiment, we initialized the penalty weight at $100$, with a growth rate of $1.05$. Table 5 presents the results under this approach. In general, we observe that the method successfully improves the feasibility at the cost of objective value. It reflects the trade-off between enhancing stricter constraint satisfaction and achieving better solutions. However, for the 20000×5 Rosenbrock problem, even with progressively increasing penalties, the predictor still yields many infeasible solutions.
@@ Appendix F Training Time Comparison
Tables 7, 8, 9 and 10, summarize the training times (in seconds) required by each method for problems of different scales. These results offer a clear view of the computational demands as problem complexity grows.

### v3

@@ 5 Learning to Optimize MINLPs with Correction Layers
The two baseline methods isolate parts of our full pipeline, serving as ablations to clarify the role of our correction layers. In Rounding after Learning (RL), rounding is applied to the relaxed solution generated by the learning-based methods after training is complete, meaning the rounding step does not participate in training. In Rounding with STE (RS), the continuous values predicted by the first neural network $\pi_{\Theta_{1}}$ are rounded to nearest integer during training using the Straight-Through Estimator (STE), enabling gradients to pass through the rounding operator, though this rounding is not learnable. For further details, refer to Appendix C.
@@ 6.3 Multi-Dimensional Mixed-Integer Rosenbrock Problem
where $\bm{x}\in\mathbb{R}^{n}$ are continuous decision variables and $\bm{y}\in\mathbb{Z}^{n}$ are integer decision variables. The vectors $\bm{p}\in\mathbb{R}^{n}$ and $\bm{q}\in\mathbb{R}^{n}$ are fixed for each instance, while the parameters $b$ and $\bm{a}$ vary. Details on how the parameters $\bm{p},\bm{q},\bm{a},$ and $b$ were generated can be found in Section D.2. Due to the structure of the constraints, it is straightforward to always construct a trivial feasible integer solution, ensuring that the problem remains feasible across instances.
@@
RC Obj Mean $23.30$ $56.99$ $505.47$ $6245.15$ $63394.52$
Obj Median $21.22$ $50.96$ $439.79$ $6500.15$ $60144.23$
% Infeasible $2\%$ $1\%$ $1\%$ $2\%$ $45\%$
@@
6.4 Effect of Increasing Penalty Weight
In this section, we evaluate the effect of progressively increasing penalty weight on the above benchmark problems. This experiment is inspired by penalty methods commonly used in optimization, such as the Augmented Lagrangian Method (Hestenes, 1969). The weight of the quadratic penalty term is typically increased throughout the iterations. This gradual increase in penalty helps avoid numerical difficulties caused by overly strict constraints at the early stages while enhancing constraint satisfaction as the solution converges.
Table 5: Comparison of Results for RC and LT with and without Increasing Penalty Weight on Benchmark Problems. See the caption of Table 2 for details.
Problem Size Method Original Increasing Penalty
Obj Mean Obj Median % Infeasible Obj Mean Obj Median % Infeasible
Convex Quadratic 100×100 RC $-13.589$ $-13.587$ $4\%$ $-10.660$ $-10.697$ $1\%$
LT $-13.745$ $-13.747$ $7\%$ $-11.548$ $-11.617$ $0\%$
200×200 RC $-31.206$ $-31.247$ $5\%$ $-25.350$ $-25.412$ $0\%$
LT $-31.261$ $-31.320$ $1\%$ $-25.110$ $-25.297$ $0\%$
500×500 RC $-73.570$ $-73.761$ $8\%$ $-60.040$ $-60.405$ $4\%$
LT $-73.457$ $-73.414$ $8\%$ $-57.507$ $-57.590$ $1\%$
Simple Non-Convex 100×100 RC $0.974$ $0.928$ $1\%$ $1.773$ $1.718$ $0\%$
[…]
200×200 RC $1.688$ $1.688$ $0\%$ $1.662$ $1.584$ $0\%$
LT $0.514$ $0.464$ $1\%$ $2.163$ $2.060$ $0\%$
@@ 6.4 Effect of Increasing Penalty Weight
Multi-Dim Rosenbrock 200×5 RC $505.47$ $439.79$ $1\%$ $519.85$ $467.59$ $0\%$
LT $608.47$ $600.25$ $0\%$ $618.51$ $624.10$ $0\%$
2000×5 RC $6245.15$ $6500.15$ $2\%$ $7276.96$ $7460.26$ $0\%$
LT $6224.79$ $6097.23$ $3\%$ $5780.85$ $5590.29$ $0\%$
20000×5 RC $63394.52$ $60144.23$ $45\%$ $63626.35$ $58977.61$ $29\%$
LT $66890.89$ $48003.27$ $32\%$ $63734.97$ $64431.54$ $39\%$
In this experiment, we initialized the penalty weight at $100$, with a growth rate of $1.05$. Table 5 compares the results under this approach. In general, we observe that the method successfully improves the feasibility at the cost of objective value. It reflects the trade-off between enhancing stricter constraint satisfaction and achieving better solutions. However, for the 20000×5 Rosenbrock problem, even with progressively increasing penalties, the predictor still yields many infeasible solutions.
6.5 Effect of Training Sample Size
The large number of infeasible solutions observed in the 20000×5 Rosenbrock problem can primarily be attributed to significant overfitting within the model. Given that we have prior knowledge of the parameter distribution and our self-supervised learning approach does not rely on optimal solution labels, we can easily scale up the sample size to effectively mitigate overfitting.
In our experiments with the 20000×5 Rosenbrock problem, we observed significant overfitting due to the large size of the neural network. Given that we have prior knowledge of the parameter distribution and our self-supervised learning approach does not rely on optimal solution labels, we can easily scale up the sample size to effectively mitigate overfitting.
To evaluate the impact of sample size on model performance, we trained our model for 20000×5 Rosenbrock on datasets of three different sizes: 800, 8,000, and 80,000 instances. We adjusted the training epochs to 2000, 200, and 20 (with early stopping), respectively, to maintain a similar number of parameter updates across different dataset sizes. All other hyperparameters were kept consistent with previous experiments to isolate the effect of sample size.
Figure 3: Performance of RC and LT Methods on the 20000×5 Rosenbrock Problem with Varying Training Sample Sizes (800, 8000, 80000).
(a) Infeasible Ratio of RC and LT Methods on Test Set across Different Training Sample Sizes.
(b) Objective Value for Each Instance of RC and LT Methods on Test Set across Different Training Sample Sizes.
As shown in Figure 3, increasing the sample size yields significant improvements in both objective values and feasibility. With 80,000 samples for training, the infeasibility ratio was reduced to 1% on the test set, demonstrating that larger sample sizes help the model generalize better to unseen instances, particularly in extremely high-dimensional settings. These results underscore the importance of adequate sample size when training models. This strategy is especially practical with our approach due to its self-supervised nature, which bypasses the need for labeled optimal solutions.
@@ Appendix F Training Time Comparison
Table 7, 8, 9 and 10, summarize the training times (in seconds) required by each method for problems of different scales. These results offer a clear view of the computational demands as problem complexity grows.